OpenStudy (anonymous):

Given f(x) = x^2 - 8 and g(x) = x +1 / x, find (gof)(1).

OpenStudy (dls):

substitute "f(x)" instead of "x" in "g(x)"

OpenStudy (anonymous):

huh?

OpenStudy (dls):

? (g(of(x))

OpenStudy (dls):

and substitute x=1 later

OpenStudy (anonymous):

its not g of x though, its g o f, like the little circle symbol thing

OpenStudy (linknissan):

hello

OpenStudy (anonymous):

hello

ganeshie8 (ganeshie8):

(gof)(1) = g(f(1))

ganeshie8 (ganeshie8):

find f(1) first, watever the value u get, plug that in g

OpenStudy (anonymous):

i just dont get it lol. im sorry. it also says Hint: (g o f)(-4) = g(f(-4))]

ganeshie8 (ganeshie8):

yes thats an useful hint :)

ganeshie8 (ganeshie8):

(g o f)(-4) g(f(-4)) g of f of (-4)

ganeshie8 (ganeshie8):

all 3 above are same, let me show u the solution, then it may become simple to see :)

OpenStudy (anonymous):

okay, thank you for taking time to help me

ganeshie8 (ganeshie8):

(gof)(1) g( f(1) ) g( 1^2-8 ) g(1-8) g(-7)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yeah i get that far

ganeshie8 (ganeshie8):

good :) rest should be easy for u. simply put x=-7 in g(x)

ganeshie8 (ganeshie8):

g(-7) -7 + 1 / -7 = ?

OpenStudy (anonymous):

-6/-7 which is simplified to 6/7?

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

thank you!!

ganeshie8 (ganeshie8):

np :)